OpenStudy (anonymous):

Tarzan is often portrayed as swinging through the jungle from vine to vine. Each of Tarzan's vines are 30 meters long and Tarzan catches each vine when it is at rest 15 degrees to the left of the (downwards) vertical, swings on it until it stops at an angle of 15 degrees to the right of vertical, and then grabs the next vine at rest and repeats the process. What is Tarzan's average horizontal speed in m/s? Details and assumptions You may take g to be 9.8 m/s2. You may treat Tarzan as a simple pendulum and use the small angle approximation.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

use the simple pendulum eq for period and velocity. \(T=2 \pi \sqrt{\frac{L}{g}}\) and \(v=\frac{2 \pi}{T}\)